Shift report 20000830 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 30 2000 - 00:22:16 EDT

  • Next message: Bjorn H Samset: "Convention?"

    ------------------------------------------------------------
    Date: Wed Aug 30 00:22:16 2000, Shift: 16:00-24:00
    Supervisor: FV, with: FV,KH
    Log book#5 pages 44 - 48
    
    Shift Summary:
    ------------------------------------------------------------
    Beam during shift from 16.00 to 20.30 where MCR called and announched
    they would debunch beam and perform studies for about an hour.
    Beam was dumped at about 21.30. Beam at injection at end of shift
    MCR needed an access to 12 o'clock.
    At about 18.00 Phobos requested dump of store, was called for
    advice. Since beam rate still quite good (75  Y.B.RC clock /sec)
    we did not recommend this. If we think beam is stil useful
    we should preferably continue.
    
    Two runs collected 
     2470  356K trigger-1 118K (6)  estimated coll. 164K i.e. 42%
     2471  282K      -     95K      estimated coll  137K 
    Both runs 90/5 deg and B polarity.
    D3D4 at 0 field.
    
    During Run 2470 trigger 4 was not operational - The rhic yellow sync
    had changed about 45 nsec. Corrected before run 2471. Trigger 4
    downscaled by 4.
    
    T3,T5 were included in data runs.
    
    
    Other accomplishments:
    a)
    - The new panels in Brahvo for T3 was tested and certified.
    - Do not use the T1,T2, MTPC1 and MTPC2 panels until checked.
    b)
      The TPC monitoring moved to SuperMonTPC to allow for higher collection
    rate for all other detectors.
      T3 monitoring added to SuperMon.
    
    Problems:
      The supermon summary plots for Tof detectors are not very useful
    e.g. the tof diff dist1  and " Time average" are essentially black
    due to dot size. 
    
     
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 30 2000 - 00:25:44 EDT